# SGS-SLAM: Semantic Gaussian Splatting for Neural Dense SLAM

### 25 Feb 2024 v2

@@ Abstract
Semantic understanding plays a crucial role in Dense Simultaneous Localization and Mapping (SLAM). Recent advancements that integrate Gaussian Splatting into SLAM systems have demonstrated its effectiveness in generating high-quality renderings. Building on this progress, we propose SGS-SLAM which provides precise 3D semantic segmentation alongside high-fidelity reconstructions. Specifically, we propose to employ multi-channel optimization during the mapping process, integrating appearance, geometric, and semantic constraints with key-frame optimization to enhance reconstruction quality. Extensive experiments demonstrate that SGS-SLAM delivers state-of-the-art performance in camera pose estimation, map reconstruction, and semantic segmentation. It outperforms existing methods by a large margin meanwhile preserves real-time rendering ability.
@@
Figure 1: The illustration of the proposed SGS-SLAM. It employs multi-channel 2D inputs encompassing appearance, geometry, and semantic information, leveraging Gaussian Splatting and differentiable rendering for joint parameter optimization. SGS-SLAM delivers accurate 3D semantic mapping along with high-fidelity map reconstructions.
**footnotetext: These authors contributed equally to this work
@@ 1 Introduction
However, NeRF-based SLAM methods employ multi-layer perceptrons (MLPs) as the implicit neural representation of scenes, which introduces several challenging limitations. Primarily, MLP models struggle with over-smoothing issues at the edge of objects, leading to a lack of fine-grained details in the map. This challenge also brings difficulties on disentangling the representation of objects, making it non-trivial to segment, edit, and manipulate objects within the scene. Moreover, when applied to larger scenes, MLP models are prone to catastrophic forgetting. This means that incorporating new scenes can adversely affect the precision of previously learned models, thereby reducing overall performance. Additionally, NeRF-based methods are computationally inefficient. Since the entire scene is modeled through one or several MLPs, it necessitates extensive model tuning for adding or updating scenes.
@@
Benefits from its rasterization of 3D primitives, Gaussian Splatting exhibits remarkably fast rendering speeds and allows direct gradient flow to each Gaussian’s parameters. This results in an almost linear projection between the dense photometric loss and parameters during optimization, unlike the hierarchical pixel-sampling and indirect gradient flow through multiple non-linear layers seen in NeRF models. Moreover, the direct projection capability simplifies the addition of new parameters to the Gaussian field as separate channels, thereby enabling dynamic multi-channel rendering. Crucially, we integrate a semantic map into the 3D Gaussian field, essential for applications in robotics and mixed-reality. This integration allows real-time rendering of appearance, depth, and semantic color.
When compared with neural implicit semantic SLAM systems, such as DNS-SLAM Li et al. (2023a) and SNI-SLAM Zhu et al. (2023), our system demonstrates remarkable superiority in terms of rendering speed, reconstruction quality, and segmentation accuracy. Leveraging these benefits, our method enable precise editing and manipulation of specific scene elements while preserving the high fidelity of the overall rendering. Furthermore, the utilization of explicit spatial and semantic information for identifying scene content can be instrumental in optimizing camera tracking. Particularly, we incorporate a two-level adjustment based on geometric and semantic criteria for key-frames selection. This process relies on recognizing objects that have been previously seen in the trajectory. Extensive experiments are conducted on both synthetic and real-world scene benchmarks. These experiments compare our method against both implicit NeRF-based approaches Yang et al. (2022); Zhu et al. (2022); Wang et al. (2023); Johari et al. (2023), and novel 3D-Gaussian-based methods Keetha et al. (2023), evaluating performance in mapping, tracking, and semantic segmentation.
@@
  • We introduce SGS-SLAM, the first semantic dense visual SLAM system grounded in 3D Gaussians. SGS-SLAM employs an explicit volumetric representation, enabling swift and real-time camera tracking and scene mapping. More importantly, it utilizes 2D semantic maps to learn 3D semantic representations expressed by Gaussians. Compared with previous NeRF-based methods which offers over-smooth object edges, SGS-SLAM provides high-fidelity reconstruction and optimal segmentation precision.
  • In SGS-SLAM, semantic maps provide additional supervision for optimizing parameters and selecting key frames. We employ a multi-channel parameter optimization strategy where appearance, geometric, and semantic signals collectively contribute to camera tracking and scene reconstruction. Furthermore, SGS-SLAM utilizes these diverse channels for the key-frame selection during the tracking phase, concentrating on actively recognizing objects seen earlier in the trajectory. This approach results in efficient and high-quality map reconstruction based on the chosen key frames.
  • Utilizing semantic representation, SGS-SLAM provides a highly accurate disentangled object representation in 3D scenes, laying a solid foundation for downstream tasks such as scene editing and manipulation. SGS-SLAM facilitates the dynamic moving, rotating, or removal of objects in the map in real-time. This is achieved by grouping Gaussians by specifying the semantic labels of the objects.
@@ 2.1 Semantic SLAM
Semantic information is of great importance for SLAM systems Mur-Artal et al. (2015); He et al. (2023); Whelan et al. (2015); Qin et al. (2018), which is a crucial requirement for applications in robotics and VR or AR fields. Real-time dense semantic SLAM systems Salas-Moreno et al. (2013); Bloesch et al. (2018); Rosinol et al. (2020) can integrate semantic information into 3D geometric representations. Traditional semantic SLAM systems rely on explicit 3D semantic expressions, such as voxel Hermans et al. (2014), point cloud Narita et al. (2019), and signed distance field Narita et al. (2019). These methods have limitations in terms of reconstruction speed, high-fidelity model acquisition, and memory usage. Moreover, traditional methods cannot reasonably infer unknown areas.
@@ 2.2 Neural Implicit SLAM
Methods based on NeRF McCormac et al. (2018), which handle complex topological structures and differentiable scene representation methods, have garnered significant attention, leading to the development of neural implicit SLAM methods Chung et al. (2023); Li et al. (2023b, 2024); Zhou et al. (2024). iMAP Sucar et al. (2021) uses a single MLP for scene representation, which shows limitation in large-scale scenes. NICE-SLAM Zhu et al. (2022) uses pre-trained multiple MLPs for hierarchical scene representation. Co-SLAM Wang et al. (2023) combines pixel set-based key-frame tracking with one-blob encoding. Go-SLAM Zhang et al. (2023) uses Droid-SLAM as front-end tracking system and multi-resolution hash encoding Müller et al. (2022) for mapping meanwhile integrated loop-closure detection and global optimization. However, these methods cannot utilize semantic information in maps. NIDS-SLAM Haghighi et al. (2023) leverages the mature front-end tracking of ORB-SLAM3 Campos et al. (2021) and Instant-NGP Müller et al. (2022) for mapping but does not optimize joint semantic features for 3D reconstruction. DNS-SLAM Li et al. (2023a) proposes a 2D semantic prior system that provides multi-view geometry constraints but does not optimize 3D reconstruction with semantic features. SNI-SLAM Zhu et al. (2023), a work parallel to ours, introduces semantic loss for geometric supervision but remains limited by the efficiency constraints of NeRF’s volume rendering.
2.3 3D Gaussian Splatting-SLAM
The outstanding performance and fast rasterization capabilities of 3D Gaussian Splatting Kerbl et al. (2023) enable higher efficiency and accuracy on sparse pixel bases. However, existing 3DGS-based SLAM systems Yan et al. (2023); Keetha et al. (2023); Huang et al. (2023) lack traditional effective tracking optimization, limiting tracking accuracy and the ability to recognize semantic information in scenes. We fuse semantic features into geometry and appearance and integrate 3D semantic features during the key-frame selection in the tracking process. This allows us to obtain more effective and higher-resolution scene segmentation results while maintaining real-time performance.
@@ 3 Method
SGS-SLAM is a Gaussian-based semantic visual SLAM systems. section 3.1 introduces its multi-channel Gaussian representation for joint parameter optimization. Like previous SLAM techniques, our method can be split into two processes: tracking and mapping. Tracking process estimates camera pose of each frame while keeping the scene parameters fixed. Mapping optimizes the scene representations based on the estimated camera pose. section 3.2 explains the breakdown steps in details. In addition, section 3.3 presents scene manipulation as a case study for downstream tasks. fig. 1 shows an overview of our system.
@@ 3.1 Multi-Channel Gaussian Representation
In order to optimize the parameters of Gaussians to represent the scene, we need to render the Gaussians into 2D images in a differentiable manner. We use the render method from Luiten et al. (2024), providing extended functionality of rendering depth in colors. It works by splatting 3D Gaussians into the image plane via approximating the integral projection of the influence function $f(\cdot)$ along the depth dimension in pixel coordinate. The center of the Gaussian $\mu$, radius $r$, and depth $d$ (in camera coordinates) is splatted using the standard point rendering formula:
@@
where $d_{i}$ denotes the depth of each Gaussian. By setting $d_{i}=1$, we are able to calculate a silhouette, $Sil_{\rm pix}=D_{\rm pix}(d_{i}=1)$, which assists in determining whether a pixel is visible in the current view. This aspect of visibility is essential for camera pose estimation, as it relies on the current reconstructed map. Additionally, it is also employed in map reconstruction, where new Gaussians are introduced in pixels lacking sufficient information.
While acquiring 3D semantic information is challenging and usually demands extensive manual labeling, 2D semantic label is a more accessible prior. In our approach, we leverage 2D semantic labels, which are often provided in datasets or can be easily obtained using state-of-the-art methods. We assign distinct channels to the parameters of Gaussians to denote their semantic labels and colors. During the rendering process, the 2D semantic map can be rendered from the reconstructed 3D scene as follows:
@@ 3.2.1 Camera Pose Estimation
Given the first frame, the camera pose is set to identity and use as the reference coordinates for the following tracking and mapping procedure. While assessing the camera pose of an RGB-D view at a new timestep, the initial camera pose is determined by adding a displacement to the previous pose, assuming constant velocity, as $E_{t+1}=E_{t}+(E_{t}-E_{t-1})$. Following this, the current pose is iteratively refined by minimizing the tracking loss between the ground truth color ($C_{\rm pix}^{GT}$), depth images ($D_{\rm pix}^{GT}$), and semantic map ($S_{\rm pix}^{GT}$) and their differentiably rendered views:
@@
3.2.2 Key-frames Selection and Weighting
@@
It represents the proportion of Gaussians captured within the camera view of the key frames. $W$ and $H$ are the width and height of the camera view. The candidates with $\eta$ lower than certain threshold $T_{\rm geo}$ are removed. After the initial geometric-based selection, a second filter is conducted based on semantic criteria. We discard key frames whose semantic maps $S_{\rm pix}$ are identical to the current frame’s semantic map, as indicated by a high mIoU score. This threshold, denoted as $T_{\rm sem}$ intents to enhance map optimization from varying viewpoints, preferring views with low mIoU overlap. The remaining candidates are randomly sampled to serve as the selected key-frames associated with the current frame. In addition, we compute an uncertainty score for each key frame, defined as $\mathcal{U}(t)=e^{-\tau t}$, with $t$ representing the timestamp of the key frame and $\tau$ being a decay coefficient. This uncertainty score is used to weight the mapping loss $\mathcal{L}_{\rm mapping}$. The intuition behind this is that key frames with a later timestamp index carry a higher uncertainty in reconstruction due to the accumulation of camera tracking errors along the trajectory.
@@ 3.2.3 Map Reconstruction
The scene is modeled using Gaussians across three distinct channels: (1) their mean coordinates represent the geometric information of the scene, (2) their appearance colors depict the scene’s visual appearance, and (3) their semantic colors indicate the semantic labels of objects. These parameters across the channels are jointly optimized during the process of Gaussian densification and optimization, while the camera pose, ascertained from tracking, remains fixed.
@@
After densification, the parameters of the map are optimized by minimizing the mapping loss as:
@@
Compared to existing NeRF-based approaches that necessitate complex model architectures and feature fusion strategies, SGS-SLAM adopts explicit Gaussian representation for mapping. This approach allows for direct gradient flow to each parameter, resulting in high rendering speeds and optimal reconstruction quality. Compared to recent Gaussian-based methods, SGS-SLAM incorporates geometric, appearance, and semantic features for multi-channel rendering. This enables the joint optimization of parameters across different channels, remarkably enhancing the efficiency and effectiveness of both mapping and segmentation processes.
3.3 Scene Manipulation
Given that the scene is represented explicitly by Gaussians, it becomes feasible to directly edit and manipulate a targeted group of Gaussians. In our case, Gaussians groups are identified based on their semantic labels. The mapping process generates these Gaussians, as defined in eq. 1, allowing for further manipulation in the following manner:

### 02 Mar 2024 v3

@@ Abstract
We present SGS-SLAM, the first semantic visual SLAM system based on Gaussian Splatting. It incorporates appearance, geometry, and semantic features through multi-channel optimization, addressing the oversmoothing limitations of neural implicit SLAM systems in high-quality rendering, scene understanding, and object-level geometry. We introduce a unique semantic feature loss that effectively compensates for the shortcomings of traditional depth and color losses in object optimization. Through a semantic-guided keyframe selection strategy, we prevent erroneous reconstructions caused by cumulative errors. Extensive experiments demonstrate that SGS-SLAM delivers state-of-the-art performance in camera pose estimation, map reconstruction, precise semantic segmentation, and object-level geometric accuracy, while ensuring real-time rendering capabilities.
@@
Figure 1: The illustration of the proposed SGS-SLAM. It employs 2D inputs encompassing appearance, geometry, and semantic information, leveraging Gaussian Splatting and differentiable rendering for multi-channel parameter optimization. During the mapping process, SGS-SLAM maps the 2D semantic prior to the 3D scene, jointly optimizing it via the mapping loss for accurate 3D segmentation outcomes.
**footnotetext: These authors contributed equally to this workfootnotetext: Corresponding author
@@ 1 Introduction
However, NeRF-based SLAM methods employ multi-layer perceptrons (MLPs) as the implicit neural representation of scenes, which introduces several challenging limitations. Primarily, MLP models struggle with over-smoothing issues at the edge of objects, leading to a lack of fine-grained details in the map. This challenge also brings difficulties in disentangling the representation of objects, making it non-trivial to segment, edit, and manipulate objects within the scene. Moreover, when applied to larger scenes, MLP models are prone to catastrophic forgetting. This means that incorporating new scenes can adversely affect the precision of previously learned models, thereby reducing overall performance. Additionally, NeRF-based methods are computationally inefficient. Since the entire scene is modeled through one or several MLPs, it necessitates extensive model tuning for adding or updating scenes.
@@
Benefits from its rasterization of 3D primitives, Gaussian Splatting exhibits remarkably fast rendering speeds and allows direct gradient flow to each Gaussian’s parameters. This results in an almost linear projection between the dense photometric loss and parameters during optimization, unlike the hierarchical pixel sampling and indirect gradient flow through multiple non-linear layers seen in NeRF models. Moreover, the direct projection capability simplifies the addition of new parameters to the Gaussian field as separate channels, thereby enabling dynamic multi-channel rendering. Crucially, we integrate a semantic map into the 3D Gaussian field, essential for applications in robotics and mixed reality. This integration allows real-time rendering of appearance, depth, and semantic color.
When compared with neural implicit semantic SLAM systems, such as DNS-SLAM Li et al. (2023a) and SNI-SLAM Zhu et al. (2023), our system demonstrates remarkable superiority in terms of rendering speed, reconstruction quality, and segmentation accuracy. Leveraging these benefits, our method enables precise editing and manipulation of specific scene elements while preserving the high fidelity of the overall rendering. Furthermore, using explicit spatial and semantic information to identify scene content can be instrumental in optimizing camera tracking. Particularly, we incorporate a two-level adjustment based on geometric and semantic criteria for keyframes selection. This process relies on recognizing objects that have been previously seen in the trajectory. Extensive experiments are conducted on both synthetic and real-world scene benchmarks. These experiments compare our method against both implicit NeRF-based approaches Yang et al. (2022); Zhu et al. (2022); Wang et al. (2023); Johari et al. (2023), and novel 3D-Gaussian-based methods Keetha et al. (2023), evaluating performance in mapping, tracking, and semantic segmentation.
@@
  • We introduce SGS-SLAM, the first semantic dense visual SLAM system grounded in 3D Gaussians. SGS-SLAM employs an explicit volumetric representation, enabling swift and real-time camera tracking and scene mapping. More importantly, it utilizes 2D semantic maps to learn 3D semantic representations expressed by Gaussians. Compared with previous NeRF-based methods which offer over-smooth object edges, SGS-SLAM provides high-fidelity reconstruction and optimal segmentation precision.
  • In SGS-SLAM, semantic maps provide additional supervision for optimizing parameters and selecting keyframes. We employ a multi-channel parameter optimization strategy where appearance, geometric, and semantic signals collectively contribute to camera tracking and scene reconstruction. Furthermore, SGS-SLAM utilizes these diverse channels for keyframe selection during the tracking phase, concentrating on actively recognizing objects seen earlier in the trajectory. This approach results in efficient and high-quality map reconstruction based on the chosen keyframes.
  • Utilizing semantic representation, SGS-SLAM provides a highly accurate disentangled object representation in 3D scenes, laying a solid foundation for downstream tasks such as scene editing and manipulation. SGS-SLAM facilitates the dynamic moving, rotating, or removal of objects in the map in real time. This is achieved by grouping Gaussians by specifying the semantic labels of the objects.
@@ 2.1 Semantic SLAM
Semantic information is of great importance for SLAM systems Mur-Artal et al. (2015); He et al. (2023); Whelan et al. (2015); Qin et al. (2018), which is a crucial requirement for applications in robotics and VR or AR fields. Real-time dense semantic SLAM systems Salas-Moreno et al. (2013); Bloesch et al. (2018); Rosinol et al. (2020) integrate semantic information into 3D geometric representations. Traditional semantic SLAM systems rely on sparse 3D semantic expressions, such as voxel Hermans et al. (2014), point cloud Narita et al. (2019), and signed distance field Narita et al. (2019). These methods struggle with accurately interpreting complex environments due to limited semantic understanding. This results in a simplified categorization of environmental features, which may not capture the full range of objects and their relationships within a space. Moreover, these methods exhibit limitations regarding reconstruction speed, high-fidelity model acquisition, and memory usage.
@@ 2.2 Neural Implicit SLAM
Methods based on NeRF McCormac et al. (2018), which handle complex topological structures and differentiable scene representation methods, have garnered significant attention, leading to the development of neural implicit SLAM methods Chung et al. (2023); Li et al. (2023b, 2024); Zhou et al. (2024). iMAP Sucar et al. (2021) uses a single MLP for scene representation, which shows limitations in large-scale scenes. NICE-SLAM Zhu et al. (2022) uses pre-trained multiple MLPs for hierarchical scene representation. Co-SLAM Wang et al. (2023) combines pixel set-based keyframe tracking with one-blob encoding. Go-SLAM Zhang et al. (2023) uses Droid-SLAM as a front-end tracking system and multi-resolution hash encoding Müller et al. (2022) for mapping meanwhile integrated loop-closure detection and global optimization. However, these methods cannot utilize semantic information in maps. NIDS-SLAM Haghighi et al. (2023) leverages the mature front-end tracking of ORB-SLAM3 Campos et al. (2021) and Instant-NGP Müller et al. (2022) for mapping but does not optimize joint semantic features for 3D reconstruction. DNS-SLAM Li et al. (2023a) proposes a 2D semantic prior system that provides multi-view geometry constraints but does not optimize 3D reconstruction with semantic features. SNI-SLAM Zhu et al. (2023), a work parallel to ours, introduces semantic loss for geometric supervision but remains limited by the efficiency constraints of NeRF’s volume rendering.
2.3 3D Gaussian Splatting SLAM
The outstanding performance and fast rasterization capabilities of 3D Gaussian Splatting Kerbl et al. (2023) enable higher efficiency and accuracy on scene reconstruction. However, existing 3DGS-based SLAM systems Yan et al. (2023); Keetha et al. (2023); Huang et al. (2023) lack traditional effective tracking optimization, limiting tracking accuracy and the ability to recognize semantic information in scenes. We fuse semantic features into geometry and appearance and integrate 3D semantic features during keyframe selection in the tracking process. This allows us to obtain more effective and higher-resolution scene segmentation results while maintaining real-time performance.
@@ 3 Method
SGS-SLAM is a Gaussian-based semantic visual SLAM system. section 3.1 introduces its multi-channel Gaussian representation for joint parameter optimization. Like previous SLAM techniques, our method can be split into two processes: tracking and mapping. The tracking process estimates the camera pose of each frame while keeping the scene parameters fixed. Mapping optimizes the scene representations based on the estimated camera pose. section 3.2 explains the breakdown steps in detail. In addition, section 3.3 presents scene manipulation as a case study for downstream tasks. fig. 1 shows an overview of our system.
@@ 3.1 Multi-Channel Gaussian Representation
In order to optimize the parameters of Gaussians to represent the scene, we need to render the Gaussians into 2D images in a differentiable manner. We use the render method from Luiten et al. (2024), providing extended functionality of rendering depth in colors. It works by splatting 3D Gaussians into the image plane via approximating the integral projection of the influence function $f(\cdot)$ along the depth dimension in pixel coordinates. The center of the Gaussian $\mu$, radius $r$, and depth $d$ (in camera coordinates) is splatted using the standard point rendering formula:
@@
where $d_{i}$ denotes the depth of each Gaussian. By setting $d_{i}=1$, we can calculate a silhouette, $Sil_{\rm pix}=D_{\rm pix}(d_{i}=1)$, which assists in determining whether a pixel is visible in the current view. This aspect of visibility is essential for camera pose estimation, as it relies on the current reconstructed map. Additionally, it is also employed in map reconstruction, where new Gaussians are introduced in pixels lacking sufficient information.
While acquiring 3D semantic information is challenging and usually demands extensive manual labeling, the 2D semantic label is more accessible prior. In our approach, we leverage 2D semantic labels, which are often provided in datasets or can be easily obtained using off-the-shelf methods. We assign distinct channels to the parameters of Gaussians to denote their semantic labels and colors. During the rendering process, the 2D semantic map can be rendered from the reconstructed 3D scene as follows:
@@
The Gaussian representations employed in SGS-SLAM facilitate high-quality reconstructions at high rendering speed, offering exceptional accuracy in capturing complex textures and geometry with remarkable detail and efficiency. Furthermore, the integration of semantic features within our method significantly advances optimal scene interpretation and precise object-level geometry, effectively mitigating the oversmoothing issues prevalent in NeRF models.
@@ 3.2.1 Camera Pose Estimation
Given the first frame, the camera pose is set to identity and used as the reference coordinates for the following tracking and mapping procedure. While assessing the camera pose of an RGB-D view at a new timestep, the initial camera pose is determined by adding a displacement to the previous pose, assuming constant velocity, as $E_{t+1}=E_{t}+(E_{t}-E_{t-1})$. Following this, the current pose is iteratively refined by minimizing the tracking loss between the ground truth color ($C_{\rm pix}^{GT}$), depth images ($D_{\rm pix}^{GT}$), and semantic map ($S_{\rm pix}^{GT}$) and their differentiably rendered views:
@@
3.2.2 Keyframes Selection and Weighting
@@
It represents the proportion of Gaussians captured within the camera view of the keyframes. $W$ and $H$ are the width and height of the camera view. The candidates with $\eta$ lower than a certain threshold $T_{\rm geo}$ are removed. After the initial geometric-based selection, a second filter is conducted based on semantic criteria. We discard keyframes whose semantic maps $S_{\rm pix}$ are identical to the current frame’s semantic map, as indicated by a high mIoU score. This threshold, denoted as $T_{\rm sem}$ intends to enhance map optimization from varying viewpoints, preferring views with low mIoU overlap. The remaining candidates are randomly sampled to serve as the selected keyframes associated with the current frame. In addition, we compute an uncertainty score for each keyframe, defined as $\mathcal{U}(t)=e^{-\tau t}$, with $t$ representing the timestamp of the keyframe and $\tau$ being a decay coefficient. This uncertainty score is used to weight the mapping loss $\mathcal{L}_{\rm mapping}$. The intuition behind this is that keyframes with a later timestamp index carry a higher uncertainty in reconstruction due to the accumulation of camera tracking errors along the trajectory.
@@ 3.2.3 Map Reconstruction
The scene is modeled using Gaussians across three distinct channels: (1) their mean coordinates represent the geometric information of the scene, (2) their appearance colors depict the scene’s visual appearance, and (3) their semantic colors indicate the semantic labels of objects. These parameters across the channels are jointly optimized during the process of Gaussian densification and optimization, whereas the camera pose, ascertained from tracking, remains fixed.
@@
After densification, the parameters of the map are optimized by minimizing the mapping loss:
@@
Compared to existing NeRF-based approaches Zhu et al. (2022); Johari et al. (2023); Li et al. (2023a); Zhu et al. (2023) that necessitate complex model architectures and feature fusion strategies, SGS-SLAM adopts explicit Gaussian representation for mapping. This approach allows for direct gradient flow to each parameter, resulting in high rendering speeds and optimal reconstruction quality. Compared to recent Gaussian-based methods Keetha et al. (2023); Yan et al. (2023), SGS-SLAM incorporates geometric, appearance, and semantic features for multi-channel rendering. This enables the joint optimization of parameters across different channels, remarkably enhancing the efficiency and effectiveness of both mapping and segmentation processes.
3.3 Scene Manipulation via Object-level Geometry
Given that the scene is represented explicitly by Gaussians, it becomes feasible to directly edit and manipulate a targeted group of Gaussians. In our case, Gaussian groups are identified based on their semantic labels. The mapping process generates these Gaussians, as defined in eq. 1, allowing for further manipulation in the following manner: